OpenStudy (redr0se03399):

What is the slope-intercept form of the linear equation 4x + 2y = 24?

OpenStudy (mhchen):

What is slope-intercept form?

OpenStudy (redr0se03399):

y=mx+b

OpenStudy (redr0se03399):

@Lowkey.S ?

OpenStudy (lowkey.s):

How are you mate?

OpenStudy (lowkey.s):

given line : 4x + 2y = 24. Convert it to Slope intercept form. 2y = -4x + 24 y = -2x + 12 Here our slope is -2 a line perpendicular to this line is one with slope m , such that m*(-2) = -1 so m = 1/2 therefore a general form of the equation that is perpendicular to 4x + 2y = 24 is y = 0.5x + b or -x + 2y = 2b where b is any real number, if we want this perpendicular line to have a common point with 4x + 2y = 24, we can choose (0, 12) b=12 then since 0 + 2* 12 = 2 * b one form is -x + 2y = 24

OpenStudy (lowkey.s):

metal if help! ;D

OpenStudy (redr0se03399):

my head hurts >.<

OpenStudy (lowkey.s):

It's Alright your gonna be fine mate! I encourage you to come back. If you need help.

OpenStudy (lowkey.s):

tag me if I'm online and you can you fan to fan?

OpenStudy (redr0se03399):

ok, so that was really good math that you did buut...

OpenStudy (lowkey.s):

I fan you and testinomal you! :D

OpenStudy (lowkey.s):

But, what..?

OpenStudy (redr0se03399):

it didn't really help me.

OpenStudy (lowkey.s):

Hmm I see mate. ._. What did you need help with?

OpenStudy (studygurl14):

Slope-intcerpt form is y=mx + b You currently have 4x + 2y = 24 Your goal is to get y by itself on the left-hand side.

OpenStudy (studygurl14):

Your first step @RedR0se03399 would be to "move" the 4x from the left side to the right side, right?

OpenStudy (redr0se03399):

@StudyGurl14 yes i know that.

OpenStudy (redr0se03399):

i know how to do it, thats not the problem here.

OpenStudy (studygurl14):

What's your problem?

OpenStudy (redr0se03399):

this is the problem.

OpenStudy (redr0se03399):

i have no idea what they are asking, how they want me to answer.

OpenStudy (studygurl14):

Basically, they just want you to use the numbers/letters at the bottom to form the correct equation in slope-intcerpt form

OpenStudy (studygurl14):

Simple drag-and-drop :)

OpenStudy (lowkey.s):

-x + 2y = 24 I said the answer...

OpenStudy (redr0se03399):

@StudyGurl14 oml, yes but HOW in what FORM do they want me to put in the answer?

OpenStudy (redr0se03399):

there are several ways to state the answer, and i can only put in one.

OpenStudy (studygurl14):

Okay, so I am assuming you know the answer right?

OpenStudy (studygurl14):

Your problem is just how to input it?

OpenStudy (redr0se03399):

@Lowkey.S there is not enough space for that answer.

OpenStudy (studygurl14):

the slope-intercept form would be y = -2x + 12 right?

OpenStudy (redr0se03399):

@StudyGurl14 yes, every way i have tried it i end up missing numbers or variables.

OpenStudy (studygurl14):

Does that help @RedR0se03399

OpenStudy (redr0se03399):

@StudyGurl14 yes very much so, thank you. i will tell you if that was the corect input answer :3

OpenStudy (studygurl14):

:)